Hi, I need to find the line passing through (1,42) and (17,314) using Gaussian Elimination. I think that is using matrices.
Answer by venugopalramana(3286)

Hi, I need to find the line passing through (1,42) and (17,314) using Gaussian Elimination. I think that is using matrices.
LET THE EQN. BE
AX+BY=1...IT IS PASSING THROUGH (1,42)..SO
1A+42B=1...........I
IT IS PASSING THROUGH (17,314)...SO...
17A+314B=1.................II
WE HAVE 2 SIMULTANEOUS EQNS. IN A AND B
SO AUGMENTED MATRIX OF THESE EQNS. IS
1...........42............1
17..........314...........1
NR2=R2-17*R1
1...............42.....................1
17-17*1=0......314-17*42=-400........=1-17*1=-16
NR2=R2/-400
1...............42.................1
0................1.................-16/-400=1/25
NR1=R1-42*R2
1................0...................1-42/25=-17/25
0................1....................1/25
HENCE A=-17/25
B=1/25
EQN.OF ST.LINE IS
-17X/25 +Y/25=1
-17X+Y=25